OpenStudy (anonymous):

The graph of y = 2x and the graph of y = 2-x are symmetrical. What is the line of symmetry?

OpenStudy (anonymous):

2-1=1 1/2=1/2 2x=2-x x=2/3 y=4/3 now u have the point and slope .

OpenStudy (anonymous):

there should be two lines of symmetry huh? and they're perpendicular to each other?

OpenStudy (anonymous):

yes and the other line slope is equal -1/(1/2))=-2 and the point didn't change